OpenStudy (anonymous):

Solve the following equation: 3x2 – 8x = 35

OpenStudy (anonymous):

(3x+7) (x-5)

OpenStudy (anonymous):

3x+7 = 0 3x = -7 x = -7/3 x - 5 = 0 x = 5 x = -7/3, 5